# VerifyThisBench: Generating Code, Specifications, and Proofs All at Once

### v1

@@ 4.4 Failure Mode Distribution
Figures 5 to 5 show clear improvements in model’s performance when partial solution templates are provided.
Specifically, partial success rates, where the verifier is able to confirm some goals, increase significantly. This suggests that templates or hints help models generate more accurate solutions. Timeout rates, where the program compiles but the verifier fails to complete within the time limit, remain relatively stable. This indicates that models are making meaningful progress toward valid proofs, and the verifier struggles to find counterexamples, implying the generated solutions are closer to correctness.
Compilation errors still dominate but tend to decrease under the relaxed setting for some models, demonstrating that context helps reduce syntax-level mistakes. However, some models like GPT4o-mini and o3-mini exhibit mixed trends, suggesting that while the template helps, the model’s internal understanding and code generation fidelity still vary.
@@ 4.5 Coherence
Table 3 reports each model’s coherence confidence, i.e. whether the model believes its generated specification matches the intended problem requirement. This metric is evaluated across generations passing the verification tool. While passing a formal verifier indicates syntactic and logical correctness, it does not address the alignment problem (i.e., whether the verified implementation perfectly algins with the user-intent expressed in natural language descriptions); hence, coherence offers complementary insight. Notably, except o3-mini and Qwen, models’s confidence is less than 50% on passed solutions.
The results reveal considerable variance across models in their self-assessment behavior. Models like o3-mini and Claude exhibit high confidence, often reporting over 80% coherence even in the zero-shot setting, suggesting strong internal certainty—though this may reflect overconfidence rather than accurate introspection. In contrast, models like GPT-4o and Llama show much more conservative estimates, with coherence below 30%, indicating either better-calibrated uncertainty or limited self-awareness. Interestingly, refinement tends to reduce overconfidence for some models (e.g., Claude) while slightly improving coherence estimation for others (e.g., GPT-4o and Deepseek), suggesting iterative attempts help align perceived and actual correctness.
We manually inspected a subset of successful solutions to validate if generated specifications align with the intended problem. Except for o3-mini, most models appear honest in their coherence self-assessments, with no false negatives found. Thus, our evaluation reflects an optimistic upper bound on model performance—assuming coherence estimates are accurate and verifier passes indicate best-case correctness. Automatically verifying the alignment between generated specifications and user intent in natural language remains an open technical challenge [24]. Tackling this alignment challenge is beyond the scope of this benchmark work and is left for future research.
Key Insights: Models show a wide range in coherence confidence level, suggesting various internal behaviors. On average, only 43% of passed solutions are judged coherent and our manual review suggests strong alignment.
@@ 4.6 Performance by Tools
VerifyThisBenchXS zero-shot - 1.35% 0.82% 0 5.22% - 0.21%
refinement - 21.47% 4.53% 1.28% 9.93% - 6.64%
Table 4 shows that all tools benefit from iterative refinement through feedback. In the VerifyThisBench setting, CBMC and Verus exhibit the most pronounced improvements, likely due to their syntactic similarity to C and Rust, making them more accessible to language models. Dafny also shows moderate gains in this setting.
In VerifyThisBenchXS, improvements are even more substantial. Dafny, in particular, demonstrates a remarkable leap—from near-zero success to over a 24% pass rate by Iteration 5, highlighting strong synergy with guided synthesis. In contrast, tools such as VeriFast, Frama-C, and Why3 remain largely stagnant on both benchmarks, suggesting either stricter syntactic or semantic constraints, or a structural mismatch with current model capabilities.
4.7 Performance by Relexation
Table 5: Overall Performance across Different Relaxation Settings
Code Specification Proof
@@ 4.7 Performance by Relexation
GPT4o 0.51% 10.26% 1.53% 8.67% 1.11% 4.44%
GPT4omini 0.51% 4.10% 1.53% 3.57% 0 2.22%
o3mini 0 7.18% 1.53% 5.10% 2.22% 7.78%
o4mini 0.51% 13.33% 2.04% 15.31% 1.11% 15.56%
Claude 1.03% 12.82% 2.55% 15.82% 0 7.78%
Gemini 1.03% 5.13% 0.51% 2.55% 1.11% 2.22%
Llama 0.51% 11.79% 1.02% 13.27% 0 5.56%
Deepseek 0.51% 16.41% 4.08% 17.86% 3.33% 11.11%
Qwen 0.51% 3.08% 1.02% 3.57% 0 0
Overall 0.57% 9.34% 1.76% 9.52% 0.99% 6.30%
Table 5 categorizes performance based on three relaxation types: specification, where the implementation and proof are given and the model fills in the spec; code, where the spec is provided and the model completes both the implementation and proof; and proof, where the full solution is available except for the loop invariant, which the model must supply. Across all categories, performance improves notably from Iter 1 (zero-shot) to Iter 5 (refinement), indicating that iterative refinement or feedback significantly aids verification success.
Among the three, specification relaxation yields the highest overall pass rates, suggesting that models are most effective when reasoning about what a program is supposed to do, given a working implementation and its proof context. Completing code implementation falls between the two, showing that models can sometimes generate plausible code. Completing loop invariant, arguably the most abstract and logically demanding task, results in lowest pass rates, though still showing solid gains with retries. This points to the inherent difficulty models face in understanding and completing partial proofs.
Key Insights: Generating the entire solution holistically (with overall pass rate of 9.34%) may not be more difficult than generating a specific one, e.g., loop invariant (with overall pass rate of 6.3%).
@@ 5 Conclusion
This study also has several limitations. Our coherence metric relies on the model’s self-assessment, which may not reliably reflect true specification alignment. Additionally, our exploration of iteration is limited to feedback-based refinement; more sophisticated approaches,such as fine-tuning on formal verification tasks or incorporating external verification guidance, may yield further insights and improvements. Future work should explore these directions to better understand and advance the formal reasoning capabilities of language models.
@@ Appendix A Composition of VerifyThisBenchXS
Table 6 presents the composition of VerifyThisBenchXS, summarizing the number of verification tasks by tool and category. It includes counts of implementations, specifications, and loop-related examples for five verification tools: Dafny, Frama-C, Vercors, Verifast, and Why3. In total, the benchmark comprises 481 tasks, with 195 implementations, 196 specifications, and 90 proof-related examples.
@@
Tool Implementaion Specification Proof Total
@@
Total 195 196 90 481
Appendix B VerifyThisBenchXS Data Source
Table 7 lists the sources of solutions used to construct VerifyThisBenchXS. It includes the year of publication, the name of the verification challenge, the verification tool used, and the authors or contributors of each solution.
@@ Appendix B VerifyThisBenchXS Data Source
Appendix C Model Versions
Appendix D Tool Versions
We report the versions of all verification tools used in our experiments to ensure reproducibility. The Verus verifier was run using version v0.2025.04.03.0f22710, while Why3 was evaluated with version v1.6.0. For Frama-C, we used version v30.0, and VeriFast experiments were conducted with version v25.02. The Dafny toolchain ran on version v4.10.0, and Vercors with v2.3.0. Finally, we used CBMC version v6.5.0.
Below we briefly describe each tool:
Appendix E Example Challenge and Solution
@@
13// Please implementat and verify that the index returned by the method max() given points to an element maximal in the array

### v2

@@ 4.4 Failure Mode Distribution
We categorize outcomes as NOGEN (no code detected), Compile Error, Timeout (compiles but exceeds verifier time budget), Partial (some but not all obligations proved), and Succeed. Figures 5 to 5 show clear improvements in model’s performance when partial solution templates are provided in the relaxed settings.
Specifically, partial success rates increase significantly, indicating that template hints help models generate more accurate solutions. Timeout rates remain relatively stable. This state indicates that models are making meaningful progress toward valid proofs, but the verifier struggles to find counterexamples on difficult obligations. Compilation errors still dominate but tend to decrease under the relaxed setting for some models, demonstrating that not needing to generate from scratch helps reduce syntax-level mistakes. However, some models like GPT4o-mini and o3-mini exhibit mixed trends, suggesting that while the template helps, the model’s internal understanding and code generation fidelity still vary.
@@ 4.5 Coherence
Table 3 reports each model’s coherence confidence, i.e. whether the model believes its generated specification matches the intended problem requirement. Importantly, this “self” alignment assessment is computed in a separate pass without chain‑of‑thought disclosure of how the answer was generated by the model and is thus a statistically independent evaluation. This metric is evaluated across the verified fraction of the outputs. While passing a formal verifier indicates syntactic and logical correctness, it does not address the alignment problem (i.e., whether the verified implementation perfectly aligns with the user-intent expressed in natural language descriptions); hence, coherence offers complementary insight. Notably, except o3-mini and Qwen, models’s confidence is less than 50% on passed solutions.
The results reveal considerable variance across models in their self-assessment behavior. Models like o3-mini and Claude exhibit high confidence, often reporting over 80% coherence even in the zero-shot setting, suggesting strong internal certainty—though this may reflect overconfidence rather than accurate introspection. In contrast, models like GPT-4o and Llama show much more conservative estimates, with coherence below 30%, indicating either better-calibrated uncertainty or limited self-awareness. Interestingly, refinement tends to reduce overconfidence for some models (e.g., Claude) while slightly improving coherence calibration for others (e.g., GPT-4o and Deepseek), suggesting iterative attempts help align perceived and actual correctness.
@@
We manually inspected a subset of successful solutions to validate if generated specifications align with the intended problem. Except for o3-mini, most models appear honest in their coherence self-assessments, with no false negatives found. Thus, our evaluation reflects an optimistic upper bound on true alignment—assuming coherence estimates are accurate and verifier passes indicate best-case correctness. Automatically verifying the alignment between generated specifications and user intent in natural language remains an open technical challenge [25]. Our benchmark serves as a valuable resource for systematically investigating this specification–intent alignment problem in future research. In addition, we explore a test-based evaluation approach, with preliminary results presented in Appendix F
Key Insights: Models show a wide range in coherence confidence level, suggesting varied internal behaviors. On average, only 43% of passed solutions are judged coherent and our manual review suggests strong alignment.
@@ 4.6 Performance by Tools
Table 4 shows that all tools benefit from iterative refinement through feedback. In the VerifyThisBench setting, CBMC and Verus exhibit the most pronounced improvements, likely due to their syntactic resemblance to C and Rust, making them more accessible to language models. Dafny also shows moderate gains in this setting. In VerifyThisBenchXS, improvements are even more substantial. Dafny, in particular, demonstrates a leap from near-zero success rate to over 21.4%; Verus observes an improvement around 10%. In contrast, tools such as VeriFast, Frama-C, and Why3 remain largely stagnant on both benchmarks, suggesting either stricter syntactic or semantic constraints, or a structural mismatch with current model capabilities.
@@
VerifyThisBenchXS zero-shot - 1.35% 0.82% 0 5.22% 7.52% 0.21%
refinement - 21.47% 4.53% 1.28% 9.93% 17.28% 6.64%
4.7 Performance by Relaxation
Table 5: Overall Performance across Different Relaxation Settings in VerifyThisBenchXS
Code Specification Loop
@@ 4.7 Performance by Relaxation
GPT4o 0.88% 11.06% 3.00% 9.87% 2.48% 6.61%
GPT4omini 0.88% 3.98% 2.15% 3.86% 1.65% 4.13%
o3mini 0.88% 7.52% 2.58% 7.72% 4.13% 10.74%
o4mini 0.88% 14.16% 5.15% 18.45% 4.96% 20.66%
Claude 2.21% 15.04% 4.29% 19.31% 4.96% 11.57%
Gemini 1.33% 6.19% 1.29% 4.72% 1.65% 6.61%
Llama 0.44% 11.95% 1.72% 12.88% 0.83% 8.26%
Deepseek 0.44% 15.49% 4.72% 19.31% 5.79% 14.05%
Qwen 1.33% 3.54% 1.29% 3.86% 1.65% 2.48%
Overall 1.05% 9.73% 2.90% 11.27% 3.20% 9.81%
Table 5 breaks down VerifyThisBenchXS results by Code‑Gen, Spec‑Gen, and Loop‑Gen. Iterative refinement consistently improves pass rates across all categories.
Among the three, spec-gen yields the highest overall pass rates, suggesting that models can more readily articulate reasoning about what a program is supposed to do, given a working implementation and its proof context. Completing loop invariant, arguably the most abstract and logically demanding task, results in pass rate lower than 10%, though still showing solid gains with retries. This points to the inherent difficulty models face in understanding and completing partial proofs.
Key Insights: Generating the entire solution holistically (overall pass rate@9.73%) may not be more difficult than generating a specific one, e.g., loop invariant (overall pass rate@9.81%).
@@ Appendix A Composition of VerifyThisBenchXS
Table 6 presents the composition of VerifyThisBenchXS, summarizing the number of verification tasks by tool and category. It includes counts of implementations, specifications, and loop-related completion tasks for six verification tools: Dafny, Frama-C, VerCors, Verifast, Why3, and Verus. In total, the benchmark comprises 580 tasks, with 226 implementations, 233 specifications, and 121 loop invariants related examples.
@@
Tool Implementaion Specification Loop Total
@@
Verus 31 27 31 99
Total 226 233 121 580
Appendix B Model Versions
Appendix C Tool Versions
We report exact toolchain versions for reproducibility and summarize each tool’s verification model. The Verus verifier was run using version v0.2025.04.03.0f22710, while Why3 was evaluated with version v1.6.0. For Frama-C, we used version v30.0, and VeriFast experiments were conducted with version v25.02. The Dafny toolchain ran on version v4.10.0, and VerCors with v2.3.0. Finally, we used CBMC version v6.5.0.
Docker container images and unified toolchain launch scripts are included in the released dataset. Below we briefly describe each tool:
Appendix D Prompt Formats
As prompt optimization was not the focus of this work, we used a simple, uniform structure for all models to ensure fairness across different tools. Each prompt consists of a system prompt describing the verification tool, followed by the problem description and task. System prompts used in our experiments are included in the released dataset (see artifact).
(1) System prompt: a concise tool description and key syntax/semantics reminders.
1You are an assistant that writes formally verified programs in <TOOL>.
2- Use <language/syntax> with pre/postconditions, assertions, and loop invariants as required.
3- The solution must compile and pass the <TOOL> verifier with a 60s timeout.
4- Do not use unsupported features: <list>.
5- Return a single <file-type>, with all annotations needed for verification.
(2) User prompt: the natural-language problem overview and the specific task.
1# Description
2<Problem overview in natural language; may include pseudo-code.>
3
4# Task
5<Explicit instruction: implement/specify/prove/refine the desired property.>
Appendix E Statistics of VerifyThisBench and VerifyThisBenchXS
Figure 6: Distributions of dataset characteristics. (Left) Word count distribution of natural language descriptions for challenges. (Right) Line-of-code distribution of collected ground-truth solutions.
To provide empirical support for our claim regarding the range of difficulty in the dataset, we report several descriptive statistics. The natural language descriptions (problem overview and task statements) vary substantially in length, with an average of 467 words, ranging from 30 to 1802 words. The distribution, shown in Figure 6 (left), indicates that most challenges fall within the 200–799 word range, with a small number extending beyond 1000 words. In terms of solution complexity, we analyzed 48 collected ground-truth implementations, which range from as short as 28 lines to 648 lines, with an average of 189.44 lines and a median of 124 lines per solution. As illustrated in Figure 6 (right), the majority of these solutions are under 300 lines, with a few extending beyond 500 lines.
Beyond length, the diversity of task types also reflects difficulty variation: 15 out of 41 challenges involve relatively simple data structures such as binary trees and one-dimensional arrays, whereas the remaining challenges address more complex structures, including linked lists, graphs, queues, and specialized task-specific data types. Additionally, 11 out of 41 challenges explicitly require memory safety proofs, further illustrating the technical depth of the dataset.
Natural language descriptions in Verina [49] have a median length of 110 and max length of 296 words, with accompanying code and specifications of up to 100 lines. Clever [41] reports proof lengths ranging from 10 to 225 lines. In contrast, our benchmark spans a much broader range of difficulty.
Appendix F Exploration of test-based specification verification
Inspired by parallel work [49], we further explore a test-based proxy to evaluate specification alignment. We manually construct desired input-output pairs of a problem, and verify them against the specifications generated by the models. We check if the inputs imply the described pre-conditions, and if the outputs satisfy the post-conditions. Our setup supports open-ended, complex verification problems, without restrictions on how the function signatures or data structures are defined. As a preliminary experiment, we evaluated on all passed or failed samples of Dafny specifications generated from the VerifyThisBench end-to-end tasks using test cases, on the following two benchmark problems:
  1. Finding the maximum in an array, and
  2. Finding the maximum in a tree.
For the array version, 87.5% of the generated specifications passed the test cases (as a reference, the model’s self-assessment on coherence: 93%). For the tree version, only 10% passed, mainly due to syntax errors, helper function verifiability, and other issues (reference on the model’s self-assessment on coherence: 87%).
These results differ from our manual evaluations and the model’s self-assessments. Model’s assessment focuses on intent alignment, whereas testing requires functional correctness. This illustrates the complementary nature of different evaluation methods.
Appendix G VerifyThisBenchXS Data Source
Table 7 lists the sources of solutions used to construct VerifyThisBenchXS. It includes the year of publication, the name of the verification challenge, the verification tool used, and the authors or contributors of each solution. We include canonical community solutions where available; in addition to the list, we contribute new Verus solutions (see Section 3.4).
@@ Appendix G VerifyThisBenchXS Data Source
Appendix H Example Challenge and Solution
@@
13// Please implement and verify that the index returned by the method max() given points to an element maximal in the array
@@
Appendix I Declaration of LLM Usage
This research evaluates LLM’s performance on formal verification tasks. As for the paper preparation, LLM is ONLY used to polish the writing.